# Mapping Dark Matter Through the Dust of the Milky Way Part I: Dust Correction and Phase Space Density

### v1

@@ Appendix A Velocity Independence of the Dust Efficiency
$\epsilon(\vec{x},\vec{v})={n_{\rm obs}(\vec{x})p_{\rm obs}(\vec{v}|\vec{x})\over n_{\rm corr}(\vec{x})p_{\rm corr}(\vec{v}|\vec{x})}$ (A.2)
@@
[…]
@@
Since the stars we consider in this work are primarily RC/RG stars that are located near the disk, they are predominately drawn from a single population – older disk stars [53]. Assuming this population is in equilibrium, then the velocity distribution at every location is expected to be the same for all stars of this population; and so we should expect $p(\vec{v}|\vec{x},M_{G})$ to be independent of $M_{G}$ and thus that the efficiency factor is velocity-independent.
@@
Support for this expectation can be found through examination of the Gaia DR3 data. We build a secondary dataset of stars within $2$ kpc that are nearby enough to be complete up to dust-corrected magnitudes of $M_{G}+\mu+A_{\rm rvs}\leq 14$, using $A_{\rm rvs}$ computed from the L22 dust map (Ref. [42]). This differs from our primary dataset, which uses the cut in Eq. (2.2) that instead relies on uncorrected magnitudes. The distribution of $M_{G}$ for stars within this dataset (both before and after correcting for dust-extinction) are shown in the left panel of Figure 10. In the right panel of this Figure, we show the mean and standard deviations of the distributions of each of the three components of the stellar velocities (in Galactocentric Cartesian coordinates), binned in $M_{G}$. Both before and after correcting for dust, it is apparent that for stars with $M_{G}\gtrsim-4$, the probability distribution of velocities is largely independent of $M_{G}$. This matches our expectation if the majority of stars are from a single population, and is the necessary condition for the efficiency factor to depend only on position.
@@ Appendix B Estimating an Efficiency from the L22 Dust Map
In this appendix, we introduce a technique for inverting an existing dust map (L22, Ref. [42]) to a reference selection function $\epsilon_{\rm L22}$ with which we can use to validate our CBE-derived selection function $\epsilon$. Note that $\epsilon_{\rm L22}$ is only used as a benchmark, and is not used in any of our downstream analyses of the estimated dust-corrected PSD or gravitational potential. L22 provides a 3-dimensional map of monochromatic extinctions $(A_{0})$ at a reference wavelength of $5500$ within our entire region of interest. The passband of the Gaia RVS is relatively narrow, so monochromatic extinctions can be converted to extinction in the RVS band ($A_{\rm rvs}$) using a simple extinction law [^4] https://www.cosmos.esa.int/web/gaia/edr3-extinction-law:
@@
where $k_{\rm rvs}=0.5385$ [54]. Note that this conversion cannot account for relative differences in extinction from structures with different absorption at $\lambda_{0}=5500$ versus the RVS band at $\lambda_{\rm RVS}\approx 8550$. This limits comparisons between our estimate of $\epsilon$ and $\epsilon_{\rm L22}$ inferred from the L22 dust map, as both trace structures in the local dust field observed in different bands. Consequently, there will be some level of inherent incompatibility between these two selection functions in some regions of space or at smaller length scales.
@@
Figure 11: Same as Figure 2, using extinction-corrected absolute magnitudes $M_{G,{\rm dust}}$ computed from the L22 dust map (Ref. [42]).
@@ Appendix C Estimating Uncertainties
In order to characterize the training, measurement, and statistical uncertainties for our models of $f_{\rm obs}$, $\Phi$, and $\epsilon$ we follow a similar procedure as in Ref. [8]. In summary, we train many different realizations of these models across different iterations of the Gaia dataset (plus errors) and network initializations. We then use the variation within the resulting $f_{\rm obs}$, $\Phi$, and $\epsilon$ networks as estimators of the various sources of error. A schematic of these error estimation techniques is shown in Figure 12.

### v2

@@ Appendix A Velocity Independence of the Dust Efficiency
$\epsilon(\vec{x},\vec{v})=\frac{n_{\rm obs}(\vec{x})p_{\rm obs}(\vec{v}|\vec{x})}{n_{\rm corr}(\vec{x})p_{\rm corr}(\vec{v}|\vec{x})}$ (A.2)
@@
$\epsilon(\vec{x},\vec{v})=\frac{\int dM_{G}\,n_{\text{obs}}(\vec{x}|M_{G})\,p_{\text{obs}}(\vec{v}|\vec{x},M_{G})\,p_{\text{obs}}(M_{G})}{\int dM_{G}\,n_{\text{corr}}(\vec{x}|M_{G})\,p_{\text{corr}}(\vec{v}|\vec{x},M_{G})\,p_{\text{corr}}(M_{G})}$ (A.4)
@@
Since the stars we consider in this work are primarily RC/RG stars that are located near the disk, they are predominately drawn from a single population – older disk stars [2021ApJ...916...88G]. Assuming this population is in equilibrium, then the velocity distribution at every location is expected to be the same for all stars of this population; and so we should expect $p(\vec{v}|\vec{x},M_{G})$ to be independent of $M_{G}$ and thus that the efficiency factor is velocity-independent.
@@
Support for this expectation can be found through examination of the Gaia DR3 data. We build a secondary dataset of stars within $2$ kpc that are nearby enough to be complete up to dust-corrected magnitudes of $M_{G}+\mu+A_{\rm rvs}\leq 14$, using $A_{\rm rvs}$ computed from the L22 dust map (Ref. [Gaia2mass2022]). This differs from our primary dataset, which uses the cut in Eq. (2.2) that instead relies on uncorrected magnitudes. The distribution of $M_{G}$ for stars within this dataset (both before and after correcting for dust-extinction) are shown in the left panel of Figure 10. In the right panel of this Figure, we show the mean and standard deviations of the distributions of each of the three components of the stellar velocities (in Galactocentric Cartesian coordinates), binned in $M_{G}$. Both before and after correcting for dust, it is apparent that for stars with $M_{G}\gtrsim-4$, the probability distribution of velocities is largely independent of $M_{G}$. This matches our expectation if the majority of stars are from a single population, and is the necessary condition for the efficiency factor to depend only on position.
@@ Appendix B Estimating an Efficiency from the L22 Dust Map
In this appendix, we introduce a technique for inverting an existing dust map (L22, Ref. [Gaia2mass2022]) to a reference selection function $\epsilon_{\rm L22}$ with which we can use to validate our CBE-derived selection function $\epsilon$. Note that $\epsilon_{\rm L22}$ is only used as a benchmark, and is not used in any of our downstream analyses of the estimated dust-corrected PSD or gravitational potential. L22 provides a 3-dimensional map of monochromatic extinctions $(A_{0})$ at a reference wavelength of $5500$ Å within our entire region of interest. The passband of the Gaia RVS is relatively narrow, so monochromatic extinctions can be converted to extinction in the RVS band ($A_{\rm rvs}$) using a simple extinction law [^4] https://www.cosmos.esa.int/web/gaia/edr3-extinction-law:
@@
where $k_{\rm rvs}=0.5385$ [2023A&A...674A...6S]. Note that this conversion cannot account for relative differences in extinction from structures with different absorption at $\lambda_{0}=5500$ Å versus the RVS band at $\lambda_{\rm RVS}\approx 8550$ Å. This limits comparisons between our estimate of $\epsilon$ and $\epsilon_{\rm L22}$ inferred from the L22 dust map, as both trace structures in the local dust field observed in different bands. Consequently, there will be some level of inherent incompatibility between these two selection functions in some regions of space or at smaller length scales.
@@
Figure 11: Same as Figure 2, using extinction-corrected absolute magnitudes $M_{G,{\rm dust}}$ computed from the L22 dust map (Ref. [Gaia2mass2022]).
@@ Appendix C Estimating Uncertainties
In order to characterize the training, measurement, and statistical uncertainties for our models of $f_{\rm obs}$, $\Phi$, and $\epsilon$ we follow a similar procedure as in Ref. [lim2023mapping]. In summary, we train many different realizations of these models across different iterations of the Gaia dataset (plus errors) and network initializations. We then use the variation within the resulting $f_{\rm obs}$, $\Phi$, and $\epsilon$ networks as estimators of the various sources of error. A schematic of these error estimation techniques is shown in Figure 12.